Solve: -x^4+200=102x^2.
Answer by stanbon(75887)

Solve: -x^4+200=102x^2.
-x^4-102x^2+200=0
x^4+102x^2-200=0
Using the quadratic formula you get:
x^2=[-102+sqrt(102^2-4(-200)]/2
x^2=[-102+sqrt11204]/2
x^2=[3.84894898]/2
x^2=1.92447449
x==1.38972543 or x=-1.38972543
OR
x^2=[-102-sqrt(102^2-4(-200)]/2
x^2=[-102-sqrt11204]/2
x^2=[-207.848990]/2
x^2=-103.92
x=10.19i or x=-10.19i
Cheers,
stan H.